CH/Doc/96/02

Agenda

Charissa at GANIL Meeting

To be held on: Tuesday, 1 October 1996, 2.00 pm

To be held at:  Birmingham

Agenda

1.  Items to be added to the agenda/Summary of past/coming experiments
2.  British Council application for travel
3.  Status of the mounts and preamps for running in SPEG
Timescale for rescheduled run
Preamps
Detector mounts
Cables
4.  Status of the 64 channel DA system
5.  Experiments for next year with Demon and MF experiment
Letter of Intent
Breakup experiment - detector requirements
Possible joint mounting with DEMON
Rebuilding of D6 room
Electronics and DA requirements
6.  MF and NMC extended stays at LPC and GANIL
7.  Nigel's visit after 9 November
8.  Status of the new grant announcement
9.  Proposed collaboration with Strasbourg for gamma studies
10. LPC budget discussions
Nigel will have the money to rebuild the DEMON support
so we will have that available in time for use in June/July '97.
Next year's budget will be requested to be 200 kF.  Requested as:
30 kF (3.9K) to build 3 beam tracking PPAC's for our expts, 70 kF (9.2K)
for DEMON + E281 expts and 100 kF (13.1K) to contribute the new array.
The latter two could be used for a SUN workstation.
11. Amplifiers and discriminators for new 256 channel array
DA system overview
Detector plans
Preamplifiers
Amplifiers - Charissa and CAEN
12. Date and agenda for technical meeting

W.N. Catford 30 Sep 1996